OpenStudy (anonymous):

what is 7xcos + 4 cos -2 sin in derivative form?

OpenStudy (helder_edwin):

do u mean find the derivative of \[ y=7x\cos x+4\cos x-2\sin x \] ?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

do u know how to compute derivatives?

OpenStudy (anonymous):

yes i am just having trouble with it this is wat i got

OpenStudy (anonymous):

-7sinx + (-4sinx) -(2cosx)

OpenStudy (helder_edwin):

i got \[\large y'=7\cos x+7x(-\sin x)+4(-\sin x)-2\cos x \]

OpenStudy (helder_edwin):

which can be rewritten as \[\large y'=5\cos x-(7x+4)\sin x \]

OpenStudy (anonymous):

Im kind of lost

OpenStudy (helder_edwin):

with which part?

OpenStudy (anonymous):

y did you put 7cosx

OpenStudy (anonymous):

with +7x

OpenStudy (helder_edwin):

u have \(7x\cos x\). to differentiate this u use the product rule: \[\large (uv)'=u'v+uv'. \] So \[\large (7x\cos x)'=(7x)'\cos x+7x(\cos x)'=7\cos x+7x(-\sin x) \]

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

thank you

OpenStudy (helder_edwin):

u r welcome